# ROLeR: Effective Reward Shaping in Offline Reinforcement Learning for Recommender Systems

### v1

@@ Abstract.
Offline Reinforcement Learning, Recommendation Systems
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, Idaho, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 4.2. World Model Learning
Reward Prediction is the core output of the world model. If we formulate the feedback of all the users towards all the items as a matrix, this matrix is usually highly sparse based on the offline interaction data, which is one typical characteristic of RS. The function of reward prediction, $\hat{r}$, is to answer the “what if” queries and complement the matrix. However, the quality of this matrix complement and its influence have not been investigated before. In recent methods (Gao et al., 2023a, b), the reward is the average of multiple world models:
@@
where $M$ is the number of world models, and $W_{j}$ is the $j$-th one.
Uncertainty Penalty is widely used in offline RL to estimate the risk of taking a certain action. One representative direction utilizes an ensemble of world models to calculate the distances within the ensemble as a measure of uncertainty. In DORL, it formulates the world model as a Gaussian probabilistic model (GPM) and calculates the uncertainty of one interaction, $\mathbf{x}$, as $P_{U}(\mathbf{x}):=\max_{k\in\mathbf{E}}\sigma^{2}_{\theta_{k}}$, where $k$ is the index in world model ensemble $\mathbf{E}$, and $\sigma^{2}_{\theta_{k}}$ is the variance of corresponding GPM. Thus, the estimated reward in Equation (5) has been changed to:
@@ 4.7. Uncertainty Penalty
(15) $\tilde{P}_{U}=\frac{1}{k}\sum_{u^{\prime}\in k\text{NN}(u)}d(u,u^{\prime}),$
where $d(\cdot)$ is the distance metric. Intuitively, it works as a complementary to our reward shaping method since it considers both the clustering quality and the conditions of the dataset to enhance the generalization. When the dataset is extremely sparse, the retrieved nearest neighbors for mutual inference may be less representative. Thus, our uncertainty penalty can effectively reduce the chance of making risky decisions. Specifically, we use the cosine distance as $d$ in our implementation. Extensive experiments about the design and estimation of uncertainty penalty are conducted in Section 6.4.
@@ 5. Performance Lower Bound
We are interested in the closeness of the estimated value function learned on $\hat{G}$ and the underlying ground truth value function corresponding to $G$ in the evaluation environment: $|V^{*}-V_{\pi}|,\forall s\in\mathbf{S}$. To conduct meaningful analysis, we need to make some mild assumptions on the distances between a user and its near neighbors. We denote $\hat{Q}^{*}$ as the optimal state-action value function of $\hat{G}$, $\mathcal{T}$ as the Bellman Operator and $d$ as the distance metric for selecting nearest neighbors.
@@
where $\epsilon$ is a little quantity and $Q_{m}$ is the maximal Q-value. This theorem provides our reward shaping method with a theoretical guarantee that when the users’ behaviors are not too irrelevant to form a meaningful clustering and the sparsity of the offline data is limited, then the policy learned with the world model is lower bounded.
@@ 6. Experiment
In this section, we begin with introducing the experiment setup including the datasets we use, evaluation metric, the state-of-the-art method, and other baselines. Then, we investigate the following questions:
@@ 6.1.1. Datasets
Dataset Usage No. of Users No. of Items Interaction Coverage
@@ 6.1.3. Baselines
$\bullet$ CIRS: Counterfactual Interactive Recommender System (Gao et al., 2023b), introduces a causal model to infer the user preference in the training of a PPO (Schulman et al., 2017) algorithm. CIRS is also a SOTA method on KuaiRec. The official repository of CIRS only covers KuaiRec in the four datasets.
@@ 6.2. Overall Performance (RQ1)
ROLeR achieves the closest single-step reward to the reference policy trained by ground truth reward functions on all datasets. The single-step reward range for KuaiRand is $0\sim 1$ and $0\sim 5$ for the other three datasets. Looking into the last row of Figure 2, the ROLeR significantly outperforms all baseline methods by a large margin on KuaiRec. This is the major contribution of more accurate reward functions. On KuaiRand, Coat, and Yahoo, though the differences of single-step reward, shown by the $R_{each}$ columns in Table 2 and 3, across all baselines are not so significant as KuaiRec, ROLeR can still utilize the ”inconspicuous” differences to steadily accumulated to significant advantages in the long-term reward.
@@ 6.5. The Impact of World Model (RQ4)
When summarizing the utility of the world model in policy learning, it trains the item embedding used as the initialization of action representation, predicts the user-item feedback as the estimated rewards, and estimates the uncertainty. In implementing ROLeR on KuaiRec, our reward model and uncertainty penalty are estimated from the offline data. We find that both initializing the action representation from a standard distribution and applying a Transformer state tracker can improve the cumulative reward in most of our testing settings as exemplified in Table 6. In this table, att and avg mean the transformer state tracker and average state tracker, respectively. $a_{G}$ and $a_{t}$ mean whether to initialize the action representation at the beginning of policy training or not. We believe the inaccuracy of the world model does not only hurt the estimation of the reward function but the action representation. In addition, as recent interaction history is available, the former Avg. state tracker used in DORL can hardly capture the order information. That is the reason we turn to the Transformer state tracker on KuaiRec. However, this is not the case for the other three datasets. The original testing matrix for KuaiRand is extremely sparse (see Table1). To make it a fully observed evaluation environment, the user-item feedback matrix is completed through emulation. We hypothesise that may explain why the order information is not so informative.
@@ 6.6. Hyperparameter Sensitivity(RQ5)
DORL 20.494 ± 2.670 ROLeR w. avg & $a_{t}$ 27.030 ± 2.560
DORL w. att 22.895 ± 1.454 ROLeR w. avg 29.273 ± 2.983
DORL w. $a_{G}$ 22.060 ± 1.098 ROLeR w. $a_{t}$ 28.874 ± 2.377

### v2

@@ Abstract.
Offline Reinforcement Learning; Recommendation Systems
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679633isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
@@ 4.2. World Model Learning
Reward Prediction is the core output of the world model. If we formulate the feedbacks of all the users towards all the items as a matrix, this matrix is usually highly sparse based on the offline interaction data, which is one typical characteristic of RS. The function of reward prediction, $\hat{r}$, is to answer the “what if” queries and complete the matrix. But the quality of this matrix complement and its influence have not been investigated. In recent methods (Gao et al., 2023a, b), the reward is the average of multiple world models
@@
where $M$ is the number of world models, and $W_{j}$ is the $j$-th one. Uncertainty Penalty is widely used in offline RL to estimate the risk of taking a certain action. One representative direction utilizes an ensemble of world models to calculate the distances within the ensemble as a measure of uncertainty. In DORL, it formulates the world model as a Gaussian probabilistic model (GPM) and calculates the uncertainty of one interaction, $\mathbf{x}$, as $P_{U}(\mathbf{x}):=\max_{k\in\mathbf{E}}\sigma^{2}_{\theta_{k}}$, where $k$ is the index in world model ensemble $\mathbf{E}$, and $\sigma^{2}_{\theta_{k}}$ is the variance of corresponding GPM. Thus, the estimated reward in Equation (5) has been changed to:
@@ 4.7. Uncertainty Penalty
(15) $\tilde{P}_{U}(u)=\frac{1}{k}\sum_{u^{\prime}\in k\text{NN}(u)}d(u,u^{\prime}),$
where $d(\cdot)$ is the distance metric and we omit $u$ for brevity. Intuitively, it works as a complementary to our reward shaping method since it considers both the clustering quality and the conditions of the dataset to enhance the generalization. When the dataset is extremely sparse, the retrieved nearest neighbors for mutual inference may be less representative. Thus, our uncertainty penalty can effectively reduce the chance of making risky decisions. Specifically, we use the cosine distance as $d$ in our implementation. Extensive experiments about the design and estimation of uncertainty penalty are conducted in Section 6.4.
@@ 5. Performance Lower Bound
We are interested in the closeness of the estimated value function learned on $\hat{G}$ and the underlying ground truth value function corresponding to $G$ in the evaluation environment: $|V^{*}-V_{\pi}|,\forall s\in\mathbf{S}$. To conduct meaningful analysis, we need to make some mild assumptions on the distances between a user and its near neighbors. We denote $\hat{Q}^{*}$ as the optimal state-action value function of $\hat{G}$, $\mathcal{T}$ as the Bellman Operator, and $d$ as the distance metric for selecting nearest neighbors.
@@
where $\epsilon$ is a little quantity and $Q_{m}$ is the maximal Q-value. This theorem provides our reward shaping method with a theoretical guarantee that when the users’ behaviors are not too irrelevant to form a meaningful clustering and the sparsity of the offline data is limited, the policy learned with the world model is lower bounded.
@@ 6. Experiment
In this section, we begin with introducing the experiment setup including the datasets, evaluation metric, the state-of-the-art method, and other baselines. Then, we investigate the following questions:
@@ 6.1.1. Datasets
Dataset Usage No. of Users No. of Items Density
@@ 6.1.2. Evaluation
All experiments are conducted an NVIDIA RTX™ A6000 with 48 GB GDDR6. Each of these requires less than 3 GPU hours, while each trial of ROLeR requires no more than 2 GPU hours.
@@ 6.1.3. Baselines
$\bullet$ CIRS: Counterfactual Interactive Recommender System (Gao et al., 2023b), introduces a causal model to infer the user preference in the training of a PPO (Schulman et al., 2017). CIRS is also a SOTA method on KuaiRec. The official repository of CIRS only covers KuaiRec in the four datasets.
@@ 6.2. Overall Performance (RQ1)
ROLeR achieves the closest single-step reward to the reference policy trained by ground truth reward functions on all datasets. The single-step reward range for KuaiRand is $0\sim 1$ and $0\sim 5$ for the other three datasets. Looking into the last row of Figure 2, the ROLeR significantly outperforms all baseline methods by a large margin on KuaiRec. This is the major contribution of more accurate reward functions. It is also evidenced by the comparison in Figure 1. On KuaiRand, Coat, and Yahoo, though the differences of single-step reward, shown by the $R_{each}$ columns in Table 2 and 3, across all baselines are not so significant as KuaiRec, ROLeR can still utilize the ”inconspicuous” differences to steadily accumulated to significant advantages in the long-term reward.
@@ 6.5. The Impact of World Model (RQ4)
When summarizing the utility of the world model in policy learning, it trains the item embedding used as the initialization of action representation, predicts the user-item feedback as the estimated rewards, and estimates the uncertainty. In implementing ROLeR on KuaiRec, our reward model and uncertainty penalty are estimated from the offline data. We find that both initializing the action representation from a standard distribution and applying a Transformer state tracker can improve the cumulative reward in most of our testing settings as exemplified in Table 6. In this table, att and avg mean the transformer state tracker and average state tracker, respectively. $a_{G}$ and $a_{t}$ mean whether to initialize the action representation at the beginning of policy training or not. We believe the inaccuracy of the world model does not only hurt the estimation of the reward function but the action representation. In addition, as recent interaction history is available, the former Avg. state tracker used in DORL can hardly capture the order information. That is the reason we turn to the Transformer state tracker on KuaiRec. However, this is not always the case for the other three datasets. The original testing matrix for KuaiRand is extremely sparse (see Table1). To make it a fully observed evaluation environment, the user-item feedback matrix is completed through emulation. We hypothesise that may explain why the order information is not so informative on the other three datasets.
@@ 6.6. Hyperparameter Sensitivity(RQ5)
DORL 20.494 ± 2.670 ROLeR w. avg & $a_{t}$ 26.006 ± 1.337
DORL w. att 22.499 ± 1.359 ROLeR w. avg 25.748 ± 1.588
DORL w. $a_{G}$ 22.060 ± 1.098 ROLeR w. $a_{t}$ 27.646 ± 2.014